OpenStudy (anonymous):

I am definitely not the best at math. Can someone please help me answer these Trinomial Factoring problems? 1. y^(2x)-y^x-20 2. a^2-5a-6 3. 10b^2+17b+3 4. 3x^2+2xy-y^2 5. x^2-2x-24 thank you so much if you can help

OpenStudy (anonymous):

put \[y^x=t,squaring \] \[\left( y^x \right)^2=t^2,\] \[y ^{2x}=t^2\] \[t^2-t-20\] make factors